OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (4, 2).

OpenStudy (amistre64):

what do you think we should try first?

OpenStudy (amistre64):

considering that your not giving anyone an real good incentive to just give you the answers.... how should we go about solving this?

OpenStudy (anonymous):

umm listen here dnt be a smart retriceokay thts why i asked if ur not gonna help the dnt type something downn(:

OpenStudy (anonymous):

have u been given the formula used to solve this type of problem?

OpenStudy (anonymous):

The formula to solve this problem, which i showed you in the previous two questions, is : y-y1=m(x-x1)

OpenStudy (anonymous):

m=slope, and (x1,y1) are your given coordinates

OpenStudy (anonymous):

slope=m=2, (x1,y1)=(4,2),,, here your x1=4 and y1=2 to find the equation of line,,,,use the formula always y-y1=m(x-x1) just plug in those numbers.. y-2 =2(x-4) now do the multiplication and subtraction

OpenStudy (anonymous):

y-2=2x-8 y =2x-8+2 y =2x-6

OpenStudy (amistre64):

you have been given pleny of good answers; and yet decide not to give any medals for them; then you proceed to ask the same kind of question and expect us to keep doling out answers for you and not get a little irritated at your demeanor? lol

OpenStudy (anonymous):

lol.....

OpenStudy (amistre64):

hey, i dont wanna hear it, just shut up and answer the questions I so graciously bestow upon you idiots and like it ... lol

OpenStudy (amistre64):

Anti-Cheating Policy What is OpenStudy's Anti-Cheating Policy? Openstudy.com is proud to be a pioneer in the online education community and is dedicated to advancing knowledge, understanding and participation in student success. Through OpenStudy.com, students, educators and subject enthusiasts from around the globe participate in a dynamic social learning experience that stands above the rest. The resources we provide, study rooms and users, are here to help you study smarter and study with friends. We believe you will actually enjoy learning now! We work diligently to prevent misuse, but ultimately, you make the choice to use or misuse the resources we provide. We require every member to abide by OpenStudy’s Anti-cheating Policy. Openstudy.com will take swift action against anyone found misusing the community’s site. In addition to refunding any membership fee, we will ban users who cheat from the site permanently. A message to students: Please understand that as a member of the OpenStudy community you agree to abide by the academic policies or honor codes of your school and course, and report violations promptly to us. Please keep in mind that Openstudy is a community for educators as well as students. A message to educators: If you believe there has been a violation of academic integrity on our site, please do not hesitate to contact us at contact@openstudy.com I agree to abide by the AntiCheating policies.